Tohban Report 2015-06-24

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 17 Jun 2015
End Date: 24 Jun 2015
Tohban: Lindsay Glesener
Tohban email:
Next Tohban: Milo Buitrago-Casas
List all reports


Solar Activity

This was a fairly active week, with AR 12371 near disk center producing many C and M class flares. Other regions near the west limb also produced flares; we should have some good occulted events coming up this week. There is still a Major Flare Watch in effect.

How many GOES flares occurred?

 Flares above B, C, M, X class were      1    34     8     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      0    30     7     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0     0     0     0

There were RHESSI flares/GOES flares 427 / 43 over the time range 17-Jun-15 24-Jun-15

Memory Management

The max SSR fill was 46%. Despite the continued replaying of some data we missed last week and the high solar activity, there were no problems with the SSR this week.

Spacecraft Status

Data Gaps

In the first half of the week there continued to be many short data gaps in the SOH and PMTRAS data, as well as an occasional one in the monitor rates, including the strangest set of gaps ever seen to this tohban: [1] That gap set occurred at the tail end of the problem in getting data through due to stale TLEs / high solar activity (=more spacecraft drag). Ops has continued to work on resolutions to these problems, which only occur during perfect storms of high solar activity and particularly bad TLEs.

Detector issues

No threshold changes this week.

D8 and D9 front livetime are low, but they have been for awhile and there have been no negative trends recently. There are no memory problems.

It was noticed that the D9 front slow threshold is high. Does it need to be this high?

Attenuator management

Due to (1) attenuator motions not occurring ideally and (2) the desire for low-energy RHESSI data for several flares in order to look for quasi-periodic pulsations that are often seen in GOES data, it was decided that we will change the livetime percentage threshold at which the thin shutter comes in to ~90% instead of the current 94.9%. This means the shutter will wait longer before coming in (if it comes in at all). The detector spectra will likely be extremely piled up in the time before the shutter insertion. One should use this spectroscopic data with caution. We'll stay in this "high count" mode for one month and then will reevaluate after the month is over to determine whether to remain in this mode.

The tohban performed some investigation into the livetime fraction at the time of the shutter motions, and found that indeed, D1 and D7 have the best livetime, and that this livetime often sticks right around 99.2% even at the best of times. During a flare the livetime often jumps fluctuates discretely between 99.2% (the old come-out threshold) and 98.8% (the newer come-out threshold). This is the reason for the behavior we saw with the old come-out threshold (99.2% livetime/0.8% deadtime) -- the livetime never goes above that value, so the thin attenuator never comes out unless forced to. With the newer come-out threshold (98.8% livetime / 1.2% deadtime), the livetime level is often above the threshold, so the shutter almost always comes out after 4 minutes.

A few slides showing the livetime around the time of the shutter motions can be found here: [2]

Afternoon update: today (6/24) we increased the threshold used to put the thin attenuator IN (A0 to A1). We went from a value of 5.1 % to 10 % This change took effect 2015-175-22:12:56 UTC.

Spacecraft Management

We're still in Active/Vigorous decimation, and still taking events at night.

Decimation Active/Vigorous
Night time data (fronts) Taking night events
Night time data (rears) Taking night events
Require extra passes? No
Requirement for moving pointer? Some replaying
Attenuator operation Still trying to optimize motions.
Personal tools